# LS-category and topological complexity of real torus manifolds and Dold manifolds of real torus type

### 12 Jan 2024 v1

@@ 3. LS-category and topological complexity of Dold manifolds of real torus type
$\tau([x,(t_{1},\dots,t_{n})])=[x,(a_{1}t_{1},\dots,a_{n}t_{n})]$ (24)
where $x\in P$, $(t_{1},\ldots,t_{n})\in\mathbb{Z}_{2}^{n}$, and $[x,(t_{1},\dots,t_{n})]$ is the equivalence class of $(x,(t_{1},\dots,t_{n}))$ under $\sim$ in (3). One can observe that the $i$-th coordinate subgroup $1\times\dots\times 1\times\mathbb{Z}_{2}\times 1\times\dots\times 1$ of $\mathbb{Z}_{2}^{n}$ determine a nontrivial involution on $M^{n}(P,\lambda)$, say $i$-th involution. We observe that, if $(t_{1},\dots,t_{n},p)\sim(s_{1},\dots,s_{n},p)$, then $(a_{1}t_{1},\dots,a_{n}t_{n},p)\sim(a_{1}s_{1},\dots,a_{n}s_{n},p)$ as $a_{i}^{-1}t_{i}^{-1}a_{i}s_{i}=t_{i}^{-1}s_{i}$. This shows that the relation $\sim$ is preserved under any involution generated by the $\mathbb{Z}_{2}$-subgroup of $\mathbb{Z}_{2}^{n}$. The following example gives involution on some $M(P,\lambda)$ outside of these types.
@@
[…]
@@
Let $M$ be an $n$-dimensional small cover such that $M^{\left<\tau\right>}$ is path-connected for some involution $\tau$ on $M$ and the $\tau\right>$-action commutes with the $Z_{2}^{n}$-action on $M$. Then, there exist $(n+1)$ many $\tau\right>$-invariant categorical sets of $X$ and
@@
We have $M\cong M^{n}(P,\lambda)$ for a $\mathbb{Z}_{2}$-characteristic pair $(P,\lambda)$ by Proposition 2.4. We also have $n+1=\mathrm{cat}(M^{n}(P,\lambda))$ by . Therefore, using [1, Section 2], one can get $n+1=\mathrm{cat}(M^{n}(P,\lambda))\leq\mathrm{cat}_{\left<\tau\right>}(M^{n}(P,\lambda))$, since the fixed point set $M^{\tau}$ is a path-connected space.
Now we describe the $\tau$-invariant categorical cover of $M^{n}(P,\lambda)$ containing $n+1$ many sets. For any vertex $v$ of $P$, let $U_{v}:=P\setminus\cup_{v\notin F}F$ be the open subset of $P$. Let $V(P)$ be the set of vertices of $P$ and
$V_{j}=\{v\in V(P)~{}:~{}\mathfrak{f}(v)=j\},$
where $\mathfrak{f}$ is the height function of $P$. The DJ-construction of $M$ gives us $\mathfrak{q}^{-1}(U_{v})=\mathbb{Z}_{2}^{n}\times U_{v}/\sim$. Let $Y_{j}=\cup_{v\in V_{j}}\mathfrak{q}^{-1}(U_{v})$ for $j=0,\dots,n$. Since the actions of $\tau\right>$ and $\mathbb{Z}_{2}^{n}$ commute, the space $\mathfrak{q}^{-1}(V_{j})$ is $\tau$-invariant. Observe that,
@@
for any $g,h\in\mathbb{Z}_{2}^{n}$ as $v$ is a fixed point of an action of $\mathbb{Z}_{2}^{n}$ on $X$. Observe that ${P=\cup_{v\in V(P)}U_{v}}$ and $U_{v}\cap U_{w}=\emptyset$ if $\mathfrak{f}(v)=j=\mathfrak{f}(w)$. Therefore, for each $0\leq j\leq n$, one can see that
@@
Since $M^{n}(P,\lambda)^{\tau}$ is path connected, $\mathfrak{q}^{-1}({Y_{j}})$ is $\tau$-invariantly contractible in $Y$. This proves the first part of the claim.
Now, it is evident that the homotopy equivalence in (26) is $\tau$-equivariant for $j=0,\ldots,n$. Therefore, the cover $\{Y_{j}~{}:~{}0\leq j\leq n\}$ is a $\tau$-equivariant categorical cover of $X$. This gives $\mathrm{cat}_{\left<\tau\right>}(X)\leq n+1$. ∎
@@
Observe that the fixed point set $M^{\mathbb{Z}_{2}}=M_{F_{i_{1}}}\cup\dots\cup M_{F_{i_{1}}}\cup M^{\mathbb{Z}_{2}^{n}}$. Then, the left inequality follows from [2, Corollary 2.9]. Let $\mathfrak{q}\colon M\to P$ be an orbit map. Then one can observe that $\{q^{-1}(U_{v})\mid v\in V(P)\}$ forms a $\mathbb{Z}_{2}$-categorical cover of $M$, where
@@
    Let $P=I^{2}$ and consider that facets are labelled by $\{F_{1},F_{2},F_{3},F_{4}\}$. Define the characteristic function as follows $\lambda(F_{1})=(1,0)=\lambda(F_{3})$, $\lambda(F_{2})=(0,1)$ and $\lambda(F_{4})=(1,1)$. Note that $M^{n}(P,\lambda)=M$ is the Klein bottle. Let $\mathbb{Z}_{2}$-action on $M$ be given by the subgroup $\mathbb{Z}_{2}\times\{1\}$. Then one can observe that $M^{\mathbb{Z}_{2}}=M_{F_{1}}\sqcup M_{F_{3}}\cong S^{1}\sqcup S^{1}$. Therefore, from Theorem 3.3 we get $\mathrm{cat}_{\mathbb{Z}_{2}}(M)=4$.
@@
    One can generalize the above example. Let $P=P_{2m}$ be the $2m$-gon. We label facets of $P_{2m}$ by $\{F_{1},F_{2},\dots F_{2m-1},F_{2m}\}$. Then note that we can define a characteristic function $\lambda$ such that $\lambda(F_{2i-1})=(1,0)$ for $1\leq i\leq m$. Then the $\mathbb{Z}_{2}$-action on $M=M^{n}(P,\lambda)$ determined by the $\mathbb{Z}_{2}$-subgroup $\mathbb{Z}_{2}\times\{1\}$ has fixed point $M^{\mathbb{Z}_{2}}=\cup_{i=1}^{m}M_{F_{2i-1}}$. Note that $M^{\mathbb{Z}_{2}}$ contains all vertices of $P_{2m}$ and it is homeomorphic to disjoint union of $m$-many circles as $M_{F_{i}}\cong S^{1}$ for each $1\leq i\leq m$. Therefore, Theorem 3.3 gives us
@@
The following result is a consequence of [6, Corollary 5.8] and Theorem 3.2.
@@
Let $M$ be an $n$-dimensional small cover such that $M^{\left<\tau\right>}$ is path-connected for some involution $\tau$ on $M$ and the $\tau\right>$-action commutes with the $Z_{2}^{n}$-action on $M$. Then,
@@
where $\tau\colon M\to M$ be an involution. Then $D(M;p_{1},\dots,p_{r})$ is a manifold of dimension $n+p_{1}+\cdots+p_{r}$. We call this manifold a Dold manifold of real torus type. Note that we have a fibre bundle
$M\hookrightarrow{}D(M,p_{1},\dots,p_{r})\stackrel{{\scriptstyle\mathfrak{p}}}{{\longrightarrow}}P(p_{1},\dots,p_{r}),$
@@
The following result describes the mod-$2$ cohomology ring of $D(M;p_{1},\dots,p_{r})$. Let $P$ be a retractable nice manifold with corners in the sense of [28, Definition 3.2], and $M^{n}(P,\lambda)$ a real torus manifold. Then $M^{n}(P,\lambda)$ has a $\mathbb{Z}_{2}^{n}$-equivariant cell-structure. Thus, the associated $\mathbb{Z}_{2}^{n}$ acts trivially on $H^{*}(M^{n}(P,\lambda);\mathbb{Z}_{2})$. We note that any simple polytope and the polytope in Example 2.3 are retractable nice manifolds with corners.
Proposition 3.6 ([29, Proposition 4.6]).
Let $p_{1}\leq\cdots\leq p_{r}$ and $M^{n}(P;\lambda)$ be a real torus manifold over a retractable $P$. Then the cohomology ring $H^{*}(D((M,p_{1},\dots,p_{r});\mathbb{Z}_{2})$ is isomorphic as a graded $\mathbb{Z}_{2}$-algebra to
@@
It follows from [9, Theorem 2.1] that $\mathrm{cl}_{\mathbb{Z}_{2}}(P(p_{1},\dots,p_{r}))=p_{1}+r-1$. Therefore, by Proposition 3.6, we get that $\mathrm{cl}_{\mathbb{Z}_{2}}(D(M;p_{1},\dots,p_{r}))=\mathrm{cl}_{\mathbb{Z}_{2}}(M)+\mathrm{cl}_{\mathbb{Z}_{2}}(P(p_{1},\dots,p_{r}))$. Thus, the inequality in (27) follows from the cup-length lower bound on category.
Now if $M$ is small cover , then it follows from [8, Proposition 2.5] and [19, Theorem 1.2]
@@
where $q$ is the smallest integer such that $M$ is covered by $q$-many $\tau$-invariant categorical sets. From Theorem 3.2 it follows that $q=n+1$. This proves the claim. ∎
@@
Let $p_{1}\leq\cdots\leq p_{r}$ and $M$ be a real torus manifold over a retractable $P$ such that $M^{\left<\tau\right>}$ is path-connected for some involution $\tau$ on $M$ and the $\tau\right>$-action commutes with the $Z_{2}^{n}$-action on $M$. Then
@@
Moreover, $\mathrm{TC}(D(M;p_{1},\dots,p_{r}))\leq 2(n+p_{1}+r)+1$ if $M$ is an $n$-dimensional small cover.
@@
Note that, using Proposition 3.6 we have
@@
Therefore, we get the left inequality of (28). The right inequality of (28) then follows from [8, Theorem 2.5]. The remaining follows from Theorem 3.2. ∎
@@
Let $M$ be a real torus manifold satisfying the hypotheses in Theorem 2.7. Then,
@@
In particular, if $n=2^{s}$ and $p_{1}=2^{t}$, then $12^{s+1}+2^{t+1}-1\leq\mathrm{TC}(D(M;p_{1}))\leq 2^{s+1}+2^{t+1}+$ 1 .
@@
If $p_{1}$ is the power of $2$, then $32p_{1}+r+\mathrm{zcl}_{\mathbb{Z}_{2}}(S^{n})\leq\mathrm{TC}(D(M;p_{1},\dots,p_{r}))\leq 2p_{1}+2r+$ 3.
4. Symmetric topological complexity of small covers
In [17], Farber and Grant studied the symmetric analogue of the motion planning problem and introduced the notion of symmetric topological complexity. In this section, we recall the basic terminologies related to symmetric topological complexity and state some results that we are going to use further. We compute the symmetric topological complexity of some classes of small covers.
Let $Y$ be a path-connected space and consider the corresponding ordered configuration space of $2$ points
$F(Y;2)=\{(x,y)\in Y\times Y\mid x\neq y\}.$
Observe that the free path space fibration ${\pi\colon PY\rightarrow Y\times Y}$ restricts to a fibration
$\pi^{\prime}\colon P^{\prime}Y\rightarrow F(Y;2),$ (30)
where $P^{\prime}Y$ is a subspace of $PY$ consisting paths with distinct endpoints. That is $P^{\prime}Y=\{\gamma\colon I\rightarrow Y~{}|~{}\gamma(0)\neq\gamma(1)\}$. Note that ${\pi^{\prime}\colon P^{\prime}Y\rightarrow F(Y;2)}$ is a $\mathbb{Z}_{2}$-equivariant map, where the group $\mathbb{Z}_{2}$ acts freely on $F(Y;2)$ via switching the coordinates, and on $P^{\prime}Y$ by sending a path $\gamma$ to its reverse path $\bar{\gamma}$. Therefore, we have the following fibration
$\pi^{\prime\prime}\colon P^{\prime}Y/\mathbb{Z}_{2}\rightarrow B(Y;2),$ (31)
where $B(Y;2)$ is the orbit space $F(Y;2)/\mathbb{Z}_{2}$, also called an unordered configuration space of $2$ points on $Y$. This map is also a fibration.
Definition 4.1.
The symmetric topological complexity of $Y$, $\mathrm{TC}^{S}(Y)$, is defined by
$\mathrm{TC}^{S}(Y):=1+\mathrm{secat}(\pi^{\prime\prime}).$
Note that the space $B(Y;2)$ is empty if and only if $Y$ is a single point. It was mentioned in [17] that, in this case, $\mathrm{TC}^{S}(Y)=1$ and if $Y$ contains more than one point, then $\mathrm{TC}^{S}(Y)\geq 2$. Moreover, it follows from [17, Example 7] that the symmetric topological complexity of a contractible space with more than two points is $2$.
We now describe a lower bound on the symmetric topological complexity given in [17]. Let $N_{Y}$ be the sub-ring of $H^{*}(Y)\otimes H^{*}(Y)$ spanned by the norm elements (i.e., the elements of the form $x\otimes y+y\otimes x$ with $x\neq y$). The following result follows from Corollary $9$, Proposition $10$, and Theorem $17$ in [17].
@@ 4. Symmetric topological complexity of small covers
The element $1\otimes y_{j}+y_{j}\otimes 1$ is same as $1\otimes y_{j}-y_{j}\otimes 1$ in $N_{M^{n}(P,\lambda)}$. Thus the zero-divisors-cup-length of $M^{n}(P,\lambda)$ is the same as the cup-length of $N_{M^{n}(P,\lambda)}$.
@@
In particular, if $n_{j}=2^{r_{j}-1}$ for $j=1,2,\dots,m$, then
$\mathrm{TC}^{S}(M^{n}(P,\lambda))\geq 2n-m+2.$
@@
This follows from Theorem 2.13, Proposition 4.2, and Remark 4.3. ∎
@@
Let $M^{n}(P,\lambda)$ be a real torus manifold satisfying the hypotheses in Theorem 2.7. Then,
@@
This follows from Theorem 2.7, Proposition 4.2 and Remark 4.3. ∎
@@
Let $M^{n}(P,\lambda)$ be a real torus manifold satisfying the hypotheses in Theorem 2.7. Then,
$\mathrm{zcl}_{\mathbb{Z}_{2}}(\mathbb{R}P^{n})+\mathrm{zcl}_{\mathbb{Z}_{2}}(\mathbb{R}P^{p_{1}})+2\leq\mathrm{TC}^{S}(D(M(P,\lambda);p_{1}))\leq 2(n+p_{1})+1.$
In particular, if $n=2^{s}$ and $p_{1}=2^{t}$, then $12^{s+1}+2^{t+1}\leq\mathrm{TC}^{S}(M^{n}(P,\lambda))\leq 2^{s+1}+2^{t+1}+$ 1 .
@@
This follows from Theorem 2.7, Proposition 4.2 and Remark 4.3. ∎
Acknowledgement. The authors would like to thank Bikramaditya Naskar and Subhankar Sau for some helpful discussion. Koushik Brahma thanks the Indian Statistical Institute Kolkata and Chennai Mathematical Institute for financial support. Navnath Daundkar thanks NBHM for the support through the grant 0204/10/(16)/2023/R&D-II/2789.

### 04 Apr 2026 v2

@@ 3. LS-category and topological complexity of Dold manifolds of real torus type
$\tau([x,(t_{1},\dots,t_{n})])=[x,(a_{1}t_{1},\dots,a_{n}t_{n})],$ (12)
where $x\in P$, $(t_{1},\ldots,t_{n})\in\mathbb{Z}_{2}^{n}$, and $x,(t_{1},\dots,t_{n})]$ is the equivalence class of the point $(x,(t_{1},\dots,t_{n}))$ under $\sim$ in (3). One can observe that the $i$-th coordinate subgroup $1\times\dots\times 1\times\mathbb{Z}_{2}\times 1\times\dots\times 1$ of $\mathbb{Z}_{2}^{n}$ determine a nontrivial involution on $M^{n}(P,\lambda)$, say $i$-th involution. We observe that, if $(p,t_{1},\dots,t_{n})\sim(p,s_{1},\dots,s_{n})$, then $(p,a_{1}t_{1},\dots,a_{n}t_{n})\sim(p,a_{1}s_{1},\dots,a_{n}s_{n})$ as $a_{i}^{-1}t_{i}^{-1}a_{i}s_{i}=t_{i}^{-1}s_{i}$. This shows that the relation $\sim$ is preserved under any involution generated by the $\mathbb{Z}_{2}$-subgroup of $\mathbb{Z}_{2}^{n}$. The following example gives involution on some $M(P,\lambda)$ outside of these types.
@@
Let $M(P,\lambda)$ be an $n$-dimensional real torus manifold such that there is an involution $\zeta^{\prime}$ on $P$ satisfying $\lambda(\zeta^{\prime}(F))=\lambda(F)$ for any facet $F$ of $P$. Then, $\zeta^{\prime}$ induces an involution $\zeta$ on $M(P,\lambda)$ which commutes with the $\mathbb{Z}_{2}^{n}$-action on $M(P,\lambda)$. Let $\mathfrak{q}\colon M(P,\lambda)\to P$ be the orbit map. Suppose that the set $\{\lambda(F)~|~F\in\mathcal{F}(P)~\mbox{and}~F\cap P^{\left<\zeta^{\prime}\right>}\neq 0\}$ generates $\mathbb{Z}_{2}^{n}$. Then one can show from the equivalence relation $\sim$ in (3) that $\mathfrak{q}^{-1}(P^{\left<\zeta^{\prime}\right>})$ is path-connected.
@@
Let $M$ be an $n$-dimensional small cover such that $M^{\left<\tau\right>}$ is path-connected for some involution $\tau$ on $M$ and the $\left<\tau\right$>-action commutes with the $\mathbb{Z}_{2}^{n}$-action on $M$. Then, there exist $(n+1)$ $\left<\tau\right$>-invariant categorical sets of $M$ and
@@
We have $M\cong M^{n}(P,\lambda)$ for a $\mathbb{Z}_{2}$-characteristic pair $(P,\lambda)$ by Proposition˜2.4. We also have $n+1=\mathrm{cat}(M^{n}(P,\lambda))$ by Theorem˜2.5. Therefore, using [1, Section 2], one can get $n+1=\mathrm{cat}(M^{n}(P,\lambda))\leq\mathrm{cat}_{\left<\tau\right>}(M^{n}(P,\lambda))$, since the fixed point set $M^{\tau}$ is a path-connected space.
Now we describe the $\tau$-invariant categorical cover of $M^{n}(P,\lambda)$ containing $n+1$ many sets. For any vertex $v$ of $P$, let $U_{v}:=P\setminus\cup_{v\notin F}F$ be the open subset of $P$. Let $V(P)$ be the set of vertices of $P$ and $V_{j}=\{v\in V(P)~:~\mathfrak{f}(v)=j\},$ where $\mathfrak{f}$ is the height function of $P$. The DJ-construction [10, Section 1] of $M$ gives us $\mathfrak{q}^{-1}(U_{v})=\mathbb{Z}_{2}^{n}\times U_{v}/\sim.$ Let $Y_{j}=\cup_{v\in V_{j}}\mathfrak{q}^{-1}(U_{v})$ for $j=0,\dots,n$. Since the actions of $\left<\tau\right$> and $\mathbb{Z}_{2}^{n}$ commute, the space $Y_{j}$ is $\tau$-invariant. Observe that,
@@
for any $g,h\in\mathbb{Z}_{2}^{n}$ as $v$ is a fixed point of an action of $\mathbb{Z}_{2}^{n}$ on $M$. Observe that $P=\cup_{v\in V(P)}U_{v}$} and $U_{v}\cap U_{w}=\emptyset$ if $\mathfrak{f}(v)=\mathfrak{f}(w)$. Therefore, for each $0\leq j\leq n$, one can see that
@@
Since $M^{n}(P,\lambda)^{\tau}$ is path connected, $Y_{j}$} is $\tau$-invariantly contractible in $M$. This proves the first part of the claim.
Now, it is evident that the homotopy equivalence in (14) is $\tau$-equivariant for $j=0,\ldots,n$. Therefore, the cover $\{Y_{j}~:~0\leq j\leq n\}$ is a $\tau$-equivariant categorical cover of $M$. This gives $\mathrm{cat}_{\left<\tau\right>}(M)\leq n+1$. ∎
@@
Observe that the fixed point set $M^{\mathbb{Z}_{2}}=M_{F_{i_{1}}}\cup\dots\cup M_{F_{i_{1}}}\cup M^{\mathbb{Z}_{2}^{n}}$. Then, the left inequality follows from [2, Corollary 2.9]. Let $\mathfrak{q}\colon M\to P$ be an orbit map. Then, one can observe that $\{q^{-1}(U_{v})\mid v\in V(P)\}$ forms a $\mathbb{Z}_{2}$-categorical cover of $M$, where
@@
    Let $P=I^{2}$ and consider that facets are labelled by $\{F_{1},F_{2},F_{3},F_{4}\}$. Define the characteristic function as follows $\lambda(F_{1})=(1,0)=\lambda(F_{3})$, $\lambda(F_{2})=(0,1)$ and $\lambda(F_{4})=(1,1)$. Note that $M^{n}(P,\lambda)=M$ is the Klein bottle. Let $\mathbb{Z}_{2}$-action on $M$ be given by the subgroup $\mathbb{Z}_{2}\times\{1\}$. Then one can observe that $M^{\mathbb{Z}_{2}}=M_{F_{1}}\sqcup M_{F_{3}}\cong S^{1}\sqcup S^{1}$. Therefore, from Theorem˜3.3 we get $\mathrm{cat}_{\mathbb{Z}_{2}}(M)=4$.
@@
    One can generalize the above example. Let $P=P_{2m}$ be the $2m$-gon. We label facets of $P_{2m}$ by $\{F_{1},F_{2},\dots F_{2m-1},F_{2m}\}$. Then note that we can define a characteristic function $\lambda$ such that $\lambda(F_{2i-1})=(1,0)$ for $1\leq i\leq m$. Then the $\mathbb{Z}_{2}$-action on $M=M^{n}(P,\lambda)$ determined by the $\mathbb{Z}_{2}$-subgroup $\mathbb{Z}_{2}\times\{1\}$ has fixed point $M^{\mathbb{Z}_{2}}=\cup_{i=1}^{m}M_{F_{2i-1}}$. Note that $M^{\mathbb{Z}_{2}}$ contains all vertices of $P_{2m}$ and it is homeomorphic to disjoint union of $m$-many circles as $M_{F_{i}}\cong S^{1}$ for each $1\leq i\leq m$. Therefore, Theorem˜3.3 gives us
@@
The following result is a consequence of [5, Corollary 5.8] and Theorem˜3.2.
@@
Let $M$ be an $n$-dimensional small cover such that $M^{\left<\tau\right>}$ is path-connected for some involution $\tau$ on $M$ and the $\left<\tau\right$>-action commutes with the $Z_{2}^{n}$-action on $M$. Then,
@@
where $\tau\colon M\to M$ is an involution. Then $D(M;p_{1},\dots,p_{r})$ is a manifold of dimension $n+p_{1}+\cdots+p_{r}$. We call this manifold a Dold manifold of real torus type. Note that we have a fiber bundle
$M\hookrightarrow{}D(M,p_{1},\dots,p_{r})\mathrel{{\mathop{\longrightarrow}\limits^{\mathfrak{p}}}}P(p_{1},\dots,p_{r}),$
@@
The following result describes the mod-$2$ cohomology ring of $D(M;p_{1},\dots,p_{r})$. Let $P$ be a retractable nice manifold with corners in the sense of [28, Definition 3.2], and $M^{n}(P,\lambda)$ a real torus manifold. Then $M^{n}(P,\lambda)$ has a $\mathbb{Z}_{2}^{n}$-equivariant cell-structure. Thus, the associated $\mathbb{Z}_{2}^{n}$ acts trivially on $H^{*}(M^{n}(P,\lambda);\mathbb{Z}_{2})$. We note that any simple polytope and the polytope in Example˜2.3 are retractable nice manifolds with corners.
Proposition 3.6 ([27, Proposition 4.6]).
Let $p_{1}\leq\cdots\leq p_{r}$ and $M$ be a real torus manifold over a retractable $P$. Then the cohomology ring $H^{*}(D((M,p_{1},\dots,p_{r});\mathbb{Z}_{2})$ is isomorphic as a graded $\mathbb{Z}_{2}$-algebra to
@@
It follows from [8, Theorem 2.1] that $\mathrm{cl}_{\mathbb{Z}_{2}}(P(p_{1},\dots,p_{r}))=p_{1}+r-1$. Therefore, by Proposition˜3.6, we get that $\mathrm{cl}_{\mathbb{Z}_{2}}(D(M;p_{1},\dots,p_{r}))=\mathrm{cl}_{\mathbb{Z}_{2}}(M)+\mathrm{cl}_{\mathbb{Z}_{2}}(P(p_{1},\dots,p_{r}))$. Thus, the inequality in (15) follows from the cup-length lower bound on category.
Now if $M$ is small cover , then it follows from [7, Proposition 2.5] and [18, Theorem 1.2]
@@
where $q$ is the smallest integer such that $M$ is covered by $q$-many $\tau$-invariant categorical sets. From Theorem˜3.2 it follows that $q=n+1$. This proves the claim. ∎
@@
Let $p_{1}\leq\cdots\leq p_{r}$ and $M$ be a real torus manifold over a retractable $P$ such that $M^{\left<\tau\right>}$ is path-connected for some involution $\tau$ on $M$ and the $\left<\tau\right$>-action commutes with the $Z_{2}^{n}$-action on $M$. Then
@@
Moreover,
$\mathrm{TC}(D(M;p_{1},\dots,p_{r}))\leq 2(n+p_{1}+r)+1$
if $M$ is an $n$-dimensional small cover.
@@
Note that, using Proposition˜3.6 we have
@@
Therefore, we get the left inequality of (16). The right inequality of (16) then follows from [7, Theorem 2.5]. The remaining follows from Theorem˜3.2. ∎
@@
Let $M$ be a real torus manifold satisfying the hypotheses in Theorem˜2.7. Then,
@@
if $n=2^{s}$ and $p_{1}=2^{t}$, then
$2^{s+1}+2^{t+1}-1\leq\mathrm{TC}(D(M;p_{1}))\leq 2^{s+1}+2^{t+1}+1.$
One can also improve the upper bound for $\mathrm{TC}(D(M;p_{1}))$ following the proof Theorem˜2.14.
Proposition 3.10.
Let $M$ be a small cover over $\prod_{i=1}^{m}\Delta^{n_{i}}$ such that $n_{i}\geq 1$. If $n_{i}$’s and $p_{1}$ are odd then
$\mathrm{TC}(D(M;p_{1}))\leq 2(n+p_{1})-m.$
@@
If $p_{1}$ is the power of $2$, then
$2p_{1}+r+\mathrm{zcl}_{\mathbb{Z}_{2}}(S^{n})\leq\mathrm{TC}(D(M;p_{1},\dots,p_{r}))\leq 2p_{1}+2r+3.$
We now provide some applications of our results to compute the symmetric topological complexity $\mathrm{TC}^{S}(Y)$ of several real Torus manifolds $Y$. In [15], Farber and Grant studied the symmetric analog of the motion planning problem and introduced the notion of symmetric topological complexity. Let $N_{Y}$ be the sub-ring of $H^{*}(Y)\otimes H^{*}(Y)$ spanned by the elements of the form $x\otimes y+y\otimes x$ with $x\neq y$. The following result follows from Corollary $9$, Proposition $10$, and Theorem $17$ in [15].
@@
The elements $1\otimes y_{j}+y_{j}\otimes 1$ and $1\otimes y_{j}-y_{j}\otimes 1$ in $N_{M^{n}(P,\lambda)}$ are the same, since we are considering the mod-$2$ cohomology ring of real torus manifolds. Consequently, the zero-divisors-cup-length of $M^{n}(P,\lambda)$ coincides with the cup-length of $N_{M^{n}(P,\lambda)}$.
@@
In particular, if $n_{j}=2^{r_{j}-1}$ for $j=1,2,\dots,m$, then $\mathrm{TC}^{S}(M^{n}(P,\lambda))\geq 2n-m+2.$
@@
This follows from Theorem˜2.12, Proposition˜3.12, and Remark 3.13. ∎
@@
Let $M^{n}(P,\lambda)$ be a real torus manifold satisfying the hypotheses in Theorem˜2.7. Then,
@@
This follows from Theorem˜2.7, Proposition˜3.12 and Remark 3.13. ∎
@@
Let $M^{n}(P,\lambda)$ be a real torus manifold satisfying the hypotheses in Theorem˜2.7. Then,
$\mathrm{zcl}_{\mathbb{Z}_{2}}(\mathbb{R}P^{n})+\mathrm{zcl}_{\mathbb{Z}_{2}}(\mathbb{R}P^{p_{1}})+2\leq\mathrm{TC}^{S}(D(M^{n}(P,\lambda);p_{1}))\leq 2(n+p_{1})+1.$
In particular, if $n=2^{s}$ and $p_{1}=2^{t}$, then
$2^{s+1}+2^{t+1}\leq\mathrm{TC}^{S}(M^{n}(P,\lambda))\leq 2^{s+1}+2^{t+1}+1.$
@@
This follows from Theorem˜2.7, Proposition˜3.12 and Remark 3.13. ∎
Acknowledgement. The authors would like to thank Bikramaditya Naskar and Subhankar Sau for the helpful discussion. Koushik Brahma thanks the Indian Statistical Institute Kolkata and Chennai Mathematical Institute for their financial support. Navnath Daundkar thanks NBHM for the support through the grant 0204/10/(16)/2023/R&D-II/2789. The third author thanks SERB (Now ANRF) India for the CRG Grant (CRG/2023/000239).